OpenStudy (anonymous):

What is y-8=7/2(x+7) rewritten in General Form? Fractions throw me off. :c

OpenStudy (texaschic101):

y - 8 = 7/2(x + 7) -- distribute y - 8 = 7/2x + 49/2 -- get rid of fractions by multiplying EVERYTHING by common denominator of 2 2y - 16 = 7x + 49 -- subtract 7x from both sides 2y - 7x - 16 = 49 -- subtract 49 from both sides 2y - 7x - 16 - 49 = 0 -- simplify -7x + 2y - 65 = 0 -- x cannot be negative so multiply both sides by -1 7x - 2y + 65 = 0

OpenStudy (anonymous):

Thank you!!! @texaschic101 Fractions just ..ew :c Thank you so much for all your help today. :) <3 xo